# Systematic collapse of the accretion disc in AGN confirmed by UV photometry and broad line spectra

### v1

@@ Abstract
The SS disc models also fail to explain the variability of AGN, which can be more than an order of magnitude (Ren et al., 2022) on time-scales of months to years in the UV/optical band (e.g., Vanden Berk et al., 2004; MacLeod et al., 2010), while the typical viscous timescale of a standard disc is thousands of years (Noda & Done, 2018; Lawrence, 2018). A potential solution is provided by the X-ray reprocessing model (Clavel et al., 1992; Cackett et al., 2007), where the variable X-rays produced by the compact hot corona illuminate the disc, driving the UV/optical variations. However, intensive monitoring projects now clearly show that the observed optical fluctuations are poorly modelled by X-ray reprocessing (e.g., the compilation of Edelson et al., 2019). It is possible that X-ray reprocessing is the driver of the observed optical variability if the hot corona geometry is also changing in a complex way with X-ray flux (Kammoun et al., 2023), but it seems more likely that the observed optical variability is indeed intrinsic, unlike the predictions of the standard disc model (Hagen & Done, 2023; Hagen et al., 2024b).
@@
These limitations were circumvented by Hagen et al. (2024a) (hereafter H24). They used eROSITA X-ray selected AGN from the eFEDS field (Brunner et al., 2022; Liu et al., 2022) to unambiguously identify the AGN, together with high-quality Subaru Hyper Suprime-Cam (HSC) multi-band images (Aihara et al., 2022) allowing for the decomposition of the host galaxy from the AGN emission by Li et al. (2024). The resulting AGN spectra were binned as a function of black hole mass (using the stellar mass as a proxy) and AGN optical luminosity. Taking the single mass bin of gives a sequence of SEDs as a function of Eddington ratio, . H24 found that the resulting SEDs show a dramatic drop of the blue disc continuum for sources with below . Their modelling predicted that the largest change in the SED should be in the UV/EUV, however this waveband was not covered by their data.
@@
While our new study generally confirms the SED models of H24, it does also reveal a small discrepancy in the higher luminosity AGN. The disc models based on optically thick, warm Comptonisation give a smoothly increasing continuum flux, whereas the UV data have a generic break at Å as seen in previous work focused on the UV bandpass (Zheng et al., 1997; Telfer et al., 2002; Shull et al., 2012). This must be intrinsic rather than connected to dust in either the AGN or the host galaxy as the X-ray selection only included unobscured AGN. This likely points to the importance of atomic physics in the disc, perhaps due to UV line driven winds (Laor & Davis, 2014).
@@ 2.1 An eFEDS-HSC AGN sample
log | - | 42.52 0.09 | 43.16 0.11 | 43.31 0.08 | 43.83 0.09 | 44.41 0.04 | 44.83 0.06 | 45.28 0.03 |
SDSS-u detected / covered | 3 / 26 | 36 / 102 | 86 / 221 | 154 / 277 | 172 / 226 | 150 / 160 | 74 / 76 | 15 / 15 |
SDSS-u (Å) | - | 2500 | 2400 | 2290 | 2255 | 2222 | 2142 | 2083 |
SDSS-u log | - | 43.31 0.05 | 43.48 0.05 | 43.79 0.02 | 44.03 0.02 | 44.44 0.02 | 44.79 0.04 | 45.19 0.02 |
KiDS-u detected / covered | 6 / 11 | 27 / 49 | 76 / 118 | 133 / 165 | 117 / 118 | 85 / 85 | 34 / 34 | 9 / 9 |
KiDS-u (Å) | - | 2419 | 2362 | 2290 | 2290 | 2205 | 2127 | 2127 |
KiDS-u log | - | 43.21 0.04 | 43.34 0.03 | 43.63 0.02 | 43.94 0.02 | 44.37 0.04 | 44.70 0.03 | 45.09 0.08 |
@@ 2.5 Galactic and intrinsic extinction
Intrinsic extinction from the host galaxy and/or nucleus is more challenging to constrain, and becomes more important at the shorter wavelengths included here than in the original study of H24. All the AGN were selected from eROSITA to have low column density cm-2, derived using a Bayesian analysis (Liu et al., 2022). These typically can have large individual uncertainties, though most are consistent with zero as a lower limit. So it is possible that there is no cold gas associated with the host galaxy and/or AGN, giving a lower limit to the extinction correction. However, we conservatively assume that the best fit is associated with dusty gas. We first assume that this is from the AGN environment, which likely has a significantly lower dust-to-gas ratio ( 1% – 10%) than the Galactic ISM (e.g., Maiolino et al., 2001; Esparza-Arredondo et al., 2021; Jun et al., 2021). There are a variety of proposed AGN dust extinction laws (e.g., Czerny et al., 2004; Gaskell et al., 2004; Gaskell & Benker, 2007; Temple et al., 2021), but we use an SMC extinction curve (Gordon et al., 2003)666For sources with FUV data bluer than 1000 Å at rest frame, we linearly extrapolate the extinction curve. as it gives more extinction in the UV than other AGN extinction curves (Li, 2007). We derive for each individual object assuming log (Jun et al., 2021). We then calculate the extinction-corrected fluxes for each band and each source. Different intrinsic extinction prescriptions are explored in Appendix A, but these do not make a large change in the derived spectra.
@@ 2.6 Intrinsic Luminosity
We subtract the host galaxy contribution in the UV by fitting an S0 galaxy template spectrum (Silva et al., 1998) (part of the SWIRE library: Polletta et al., 2007) to the HSC host galaxy optical phototometry (see Figure A1 in H24). We use this to calculate the rest-frame monochromatic luminosity of the host galaxy in each UV band for each source. We then derive the stacked intrinsic AGN luminosity in each bin by subtracting the mean luminosity of the corresponding host galaxies (Table 1).
There is an apparent excess of the -band luminosity in the [42.5, 42.9] bin, which would vanish if compared to the HSC data without host decomposition, indicating that we have likely underestimated the galaxy contribution by using a S0 template. We note that using the single S0 template for the host will systematically underestimate the galaxy contribution for objects where the host is disc dominated rather than bulge dominated (Sersic index ). This is more often the case for objects in the lowest luminosity bins (Li et al., 2024), hence the absence of strong intrinsic UV emission will be more significant for these faint bins if we adopt an UV-brighter template. Meanwhile, the choice of galaxy templates will barely alter the intrinsic SED of the bright bins where the host galaxy is always negligible compared to the AGN in the UV.
@@ 3 The SEDs including UV data
The models are also in good agreement with the new data in the three moderate luminosity bins, following the predicted rise of the optically thick component. However, the two highest luminosity bins start to show a small deviation from the model predictions. The data turn down in the SED between two GALEX bands, corresponding to a break at 800 Å – 1400 Å. This has been widely observed before (Zheng et al., 1997; Telfer et al., 2002; Shull et al., 2012; Stevans et al., 2014; Lusso et al., 2015) but our sample clearly show that this is not an artefact of dust extinction as the is lower in these bins (see Table 1). This turnover appears intrinsic, indicating AGN with produce less UV emission than the prediction of a warm Comptonised disc, or a standard disc which predicts an even bluer spectrum (see e.g. Fig 1 of Kubota & Done, 2018).
@@
While these continua match very well in general to the new data, they are continuum models, and predict a continuous increase in the blue/UV extent of the warm Comptonisation region (see schematic at bottom right of Fig. 1, taken from H24). Hence they cannot match the downturn that appears in the data at fixed Å. This may instead indicate that atomic processes are also important in forming this feature, such as produced by e.g. a UV line driven wind, which can be efficiently produced for temperatures above 50,000 K (Laor & Davis, 2014). Alternatively it could be explained by different disc-based models, e.g., the magnetically-dominated accretion discs simulated by Hopkins (2024), in which the effective temperature has a much weaker dependence on mass and mass-accretion rate. Testing these will be the subject of a future paper.
@@ 4 Impact on changing UV ionising spectrum on the BLR
Figure 3 shows this even more clearly from a composite spectrum of the the lowest three luminosity bins (log in [42.1, 43.3]) compared to that of the last three bins (log in [44.1, 45.3]). We subtract a constant continuum component to illustrate the profiles of the lines (e.g., Mahmoud & Done, 2020), as shown in the inset of Figure 3. These spectra are normalised to match the narrow [O iii] luminosity, showing a clear change in the H line. The broad component dominates at high luminosities, but it is absent or extremely weak for sources with 0.01. This gives further evidence that there is an intrinsic change in UV continuum for sources with 0.01. We stress again that these all have low column densities so the BLR cannot simply be obscured.
@@ 5 Discussion and Conclusions
We have extended the sequence of optical-X-ray SEDs of AGN of H24. The resulting stacked optical/X-ray SEDs for black holes at fixed mass show a dramatic transition, similar to that seen in stellar mass black hole binaries, where the dominating disc component in bright AGN collapses into an inefficient X-ray plasma below . The models fit to these datasets in H24 predicted the largest change in SED should be in the rest frame UV, which was not covered by their data, so we critically test this by adding archival photometric data from SDSS-u, KiDS-u and GALEX-NUV and GALEX-FUV. We subtract a standard S0 host galaxy spectrum from these, and correct the AGN luminosity for SMC-like dust extinction assuming the Baysian best fit derived from the X-ray spectral fits of Liu et al. (2022). The resulting stacked spectra confirm that there is indeed an abrupt transition of the accretion structure, as claimed by H24, very similar to that observed in stellar mass black holes and a few individual ‘changing-look’ AGN.
@@ Acknowledgements
Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs 177.A-3016, 177.A-3017, 177.A-3018 and 179.A-2004, and on data products produced by the KiDS consortium. The KiDS production team acknowledges support from: Deutsche Forschungsgemeinschaft, ERC, NOVA and NWO-M grants; Target; the University of Padova, and the University Federico II (Naples).
@@
The optical and X-ray data as well as the SEDs are from Hagen et al. (2024a). The GALEX data are available at https://galex.stsci.edu/casjobs/default.aspx, SDSS data at https://skyserver.sdss.org/CasJobs/, and KiDS data at https://www.eso.org/qi/catalogQuery/index/260. The data in this paper will be shared on reasonable request to the corresponding author.

### v2

@@ Abstract
The SS disc models also fail to explain the variability of AGN, which can be more than an order of magnitude (Ren et al., 2022) on time-scales of months to years in the UV/optical band (e.g., Vanden Berk et al., 2004; MacLeod et al., 2010), while the typical viscous timescale of a standard disc is thousands of years (Noda & Done, 2018; Lawrence, 2018). A potential solution is provided by the X-ray reprocessing model (Clavel et al., 1992; Cackett et al., 2007), where the variable X-rays produced by the compact hot corona illuminate the disc, driving the UV/optical variations. However, intensive monitoring projects now clearly show that the observed optical fluctuations are poorly modelled by X-ray reprocessing (e.g., the compilation of Edelson et al., 2019). It is possible that X-ray reprocessing is the driver of the observed optical variability if the hot corona geometry is also changing in a complex way with X-ray flux (Kammoun et al., 2023), but it seems more likely that the observed optical variability is indeed intrinsic (Cai et al., 2018; Cai et al., 2020), unlike the predictions of the standard disc model (Hagen & Done, 2023; Hagen et al., 2024a).
@@
These limitations were circumvented by Hagen et al. (2024b) (hereafter H24). They used eROSITA X-ray selected AGN from the eFEDS field (Brunner et al., 2022; Liu et al., 2022) to unambiguously identify the AGN, together with high-quality Subaru Hyper Suprime-Cam (HSC) multi-band images (Aihara et al., 2022) allowing for the decomposition of the host galaxy from the AGN emission by Li et al. (2024). The resulting AGN spectra were binned as a function of black hole mass (using the stellar mass as a proxy) and AGN optical luminosity. Taking the single mass bin of gives a sequence of SEDs as a function of Eddington ratio, . H24 found that the resulting SEDs show a dramatic drop of the blue disc continuum for sources with below . Their modelling predicted that the largest change in the SED should be in the UV/EUV, however this waveband was not covered by their data.
@@
While our new study generally confirms the SED models of H24, it does also reveal a small discrepancy in the higher luminosity AGN. The disc models based on optically thick, warm Comptonisation give a smoothly increasing continuum flux, whereas the UV data have a generic break at Å as seen in previous work focused on the UV bandpass (Zheng et al., 1997; Telfer et al., 2002; Shull et al., 2012; Cai & Wang, 2023; Cai, 2024). This must be intrinsic rather than connected to dust in either the AGN or the host galaxy as the X-ray selection only included unobscured AGN. This likely points to the importance of atomic physics in the disc, perhaps due to UV line driven winds (Laor & Davis, 2014).
@@ 2.1 An eFEDS-HSC AGN sample
log | - | - | 43.06 0.09 | 43.40 0.12 | 43.93 0.09 | 44.42 0.04 | 44.80 0.05 | 45.27 0.02 |
SDSS- detected / covered | 3 / 26 | 36 / 102 | 86 / 221 | 154 / 277 | 172 / 226 | 150 / 160 | 74 / 76 | 15 / 15 |
SDSS- (Å) | - | 2500 | 2400 | 2290 | 2255 | 2222 | 2142 | 2083 |
SDSS- log | - | 43.00 0.05 | 43.30 0.05 | 43.70 0.02 | 43.97 0.02 | 44.42 0.03 | 44.78 0.04 | 45.18 0.02 |
KiDS- detected / covered | 6 / 11 | 27 / 49 | 76 / 118 | 133 / 165 | 117 / 118 | 85 / 85 | 34 / 34 | 9 / 9 |
KiDS- (Å) | - | 2419 | 2362 | 2290 | 2290 | 2205 | 2127 | 2127 |
KiDS- log | - | 42.83 0.04 | 43.06 0.02 | 43.48 0.02 | 43.87 0.02 | 44.34 0.03 | 44.68 0.03 | 45.08 0.03 |
@@
*All values are stacked luminosities with units of and all values are mean effective rest-frame wavelengths of each band. The normalized accretion rates (i.e., the Eddington ratio) were derived in H24 by fitting the HSC and eROSITA data with the model agnsed. Note hard X-ray dominates the bolometric luminosity in low accretion bins, which are not fully shown in Figure 1. is from the eFEDS AGN catalog, with unit of cm-2.
@@ 2.5 Galactic and intrinsic extinction
Intrinsic extinction from the host galaxy and/or nucleus is more challenging to constrain, and becomes more important at the shorter wavelengths included here than in the original study of H24. All the AGN were selected from eROSITA to have low column density cm-2, derived using a Bayesian analysis (Liu et al., 2022). These typically can have large individual uncertainties, though most are consistent with zero as a lower limit. So it is possible that there is no cold gas associated with the host galaxy and/or AGN, giving a lower limit to the extinction correction. However, we conservatively assume that the best fit is associated with dusty gas. We first assume that this is from the AGN environment, which likely has a significantly lower dust-to-gas ratio ( 1% – 10%) than the Galactic ISM (e.g., Maiolino et al., 2001; Esparza-Arredondo et al., 2021; Jun et al., 2021). There are a variety of proposed AGN dust extinction laws (e.g., Czerny et al., 2004; Gaskell et al., 2004; Gaskell & Benker, 2007; Temple et al., 2021), but we use an SMC extinction curve (Gordon et al., 2003)666For sources with FUV data bluer than 1000 Å at rest frame, we linearly extrapolate the extinction curve. as it gives more extinction in the UV than other AGN extinction curves (Li, 2007). If we adopt a flatter extinction curve (less extinction in the UV), the estimated intrinsical UV luminosity would decrease, making the lack of UV emission in the low accretion bins even more significant. We derive for each individual object assuming log (Jun et al., 2021). We then calculate the extinction-corrected fluxes for each band and each source. Different intrinsic extinction prescriptions are explored in Appendix A, but these do not make a large change in the derived spectra.
@@ 2.6 Intrinsic Luminosity
We subtract the host galaxy contribution in the UV by fitting an Sb galaxy template spectrum (Silva et al., 1998) (part of the SWIRE library: Polletta et al., 2007) to the HSC host galaxy optical phototometry (see Figure A1 in H24). We use this to calculate the rest-frame monochromatic luminosity of the host galaxy in each UV band for each source. We then derive the stacked intrinsic AGN luminosity in each bin by subtracting the mean luminosity of the corresponding host galaxies (Table 1).
We adopt a Sb template because the host galaxy in the majority of our sources is disc dominated rather than bulge dominated (65% have Sersic index , Li et al., 2024), and we do see lots of spiral features in the HSC fitting residual images by visual check. We confirm the absence of strong intrinsic UV emission will remain significant for these faint bins even if adopting an UV-fainter template (e.g., S0). Meanwhile, the choice of galaxy templates will barely alter the intrinsic SED of the bright bins where the host galaxy is always negligible compared to the AGN in the UV.
2.7 Variability
Our optical-to-X-ray SED consists of data from five different intruments spanning 20 years. Most SDSS photometries were conducted between 1998 and 2002, with a minority conducted in 2006; GALEX data were derived by coadding the data spanning 2003 to 2012; KiDS data were taken between 2011 and 2018, HSC photometries between 2014 and 2020, and eFEDS data in 2019. By stacking the SEDs of a large number of sources, we should have largely eliminated the effect of stochastic variations. Meanwhile, although dramatic “changing-state” variations can happen within years or decades (Noda & Done, 2018), such events should be too rare to bias our stacked SED. Furthermore, by comparing the -band photometries of SDSS and KiDS in individual sources, we confirm our sample shows no systematic brightening or darkening between 2000 and 2015, and the stacked -band data are also well consistent. We thus conclude the AGN variability should only contribute to some scatter of the stacked SED instead of biasing the shape.
@@ 3 The SEDs including UV data
The models are also in good agreement with the new data in the three moderate luminosity bins, following the predicted rise of the optically thick component. However, the three highest luminosity bins start to show a small deviation from the model predictions. The data turn down in the SED between two GALEX bands, corresponding to a break at 800 Å – 1400 Å. This has been widely observed before (Zheng et al., 1997; Telfer et al., 2002; Shull et al., 2012; Stevans et al., 2014; Lusso et al., 2015) but our sample clearly show that this is not an artefact of dust extinction as the is lower in these bins (see Table 1). This turnover appears intrinsic, indicating AGN with produce less UV emission than the prediction of a warm Comptonised disc, or a standard disc which predicts an even bluer spectrum (see e.g. Fig 1 of Kubota & Done, 2018).
@@
While these continua match very well in general to the new data, they are continuum models, and predict a continuous increase in the blue/UV extent of the warm Comptonisation region (see schematic at bottom right of Fig. 1, taken from H24). Hence they cannot match the downturn that appears in the data at fixed Å (see the Appendix C for the model re-fitted including the UV data). This may instead indicate that atomic processes are also important in forming this feature, such as produced by e.g. a UV line driven wind, which can be efficiently produced for temperatures above 50,000 K (Laor & Davis, 2014). Alternatively it could be explained by different disc-based models, e.g., the magnetically-dominated accretion discs simulated by Hopkins (2024), in which the effective temperature has a much weaker dependence on mass and mass-accretion rate. Testing these will be the subject of a future paper.
@@ 4 Impact on changing UV ionising spectrum on the BLR
Figure 3 shows this even more clearly from a composite spectrum of the the lowest three luminosity bins (log in [42.1, 43.3]) compared to that of the last three bins (log in [44.1, 45.3]). We subtract a constant continuum component to illustrate the profiles of the lines (e.g., Mahmoud & Done, 2020), as shown in the inset of Figure 3. These spectra are normalised to match the narrow [O iii] luminosity, showing a clear change in the H line. The broad component dominates at high luminosities, but it is absent or extremely weak for sources with 0.01888For a possible alternative scenario, HST narrow slit spectroscopy has revealed extremely broad emission lines in a low acrretion (log ) AGN NGC 3147(Bianchi et al., 2019; Bianchi et al., 2022). SDSS spectra cannot distinguish them.. This gives further evidence that there is an intrinsic change in UV continuum for sources with 0.01. We stress again that these all have low column densities so the BLR cannot simply be obscured.
@@ 5 Discussion and Conclusions
We have extended the sequence of optical-X-ray SEDs of AGN of H24. The resulting stacked optical/X-ray SEDs for black holes at fixed mass show a dramatic transition, similar to that seen in stellar mass black hole binaries, where the dominating disc component in bright AGN collapses into an inefficient X-ray plasma below . The models fit to these datasets in H24 predicted the largest change in SED should be in the rest frame UV, which was not covered by their data, so we critically test this by adding archival photometric data from SDSS-u, KiDS-u and GALEX-NUV and GALEX-FUV. We subtract a standard Sb host galaxy spectrum from these, and correct the AGN luminosity for SMC-like dust extinction assuming the Baysian best fit derived from the X-ray spectral fits of Liu et al. (2022). The resulting stacked spectra confirm that there is indeed an abrupt transition of the accretion structure, as claimed by H24, very similar to that observed in stellar mass black holes and a few individual ‘changing-look’ AGN.
@@ Acknowledgements
Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs 177.A-3016, 177.A-3017, 177.A-3018 and 179.A-2004, and on data products produced by the KiDS consortium. The KiDS production team acknowledges support from: Deutsche Forschungsgemeinschaft, ERC, NOVA and NWO-M grants; Target; the University of Padova, and the University Federico II (Naples). This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France.
@@
The optical and X-ray data as well as the SEDs are from Hagen et al. (2024b). The GALEX data are available at https://galex.stsci.edu/casjobs/default.aspx, SDSS data at https://skyserver.sdss.org/CasJobs/, and KiDS data at https://www.eso.org/qi/catalogQuery/index/260. The data in this paper will be shared on reasonable request to the corresponding author.
@@ Appendix B The intrinsic distribution of in each luminosity bin
Appendix C Fitting the full-band SED including the UV data
Here we present the best-fit models by fitting the full-band SED including the new UV data in Figure 6, and the main results stay the same. Some contribution from the warm corona is now needed in the second and third faintest bins due to a tiny excess of the UV data, though we stress large uncertainties exist in these photometries induced by the estimation of host contamination and intrinsic extinction. Meanwhile, for the three brightest bins, the warm corona becomes weaker and the acrretion rate slightly drops, due to the FUV break. Trying to fit this break also leads to fitting residuals in the X-ray band, especially in the second and third brightest bins.